OpenStudy (anonymous):

what is the derivative of (ln(2))(cos 1/4 - x)

OpenStudy (anonymous):

i dont understand why its not a negative number, because the derivative of cos(any x) is -sin(any x)

sam (.sam.):

(ln(2))(cos((1)/(4)-x)) cos((1)/(4)-x) (d)/(dx) 1/4-x=-1 d/dx =sin((1/(4)-x)) sin(1/(4)-x)

OpenStudy (anonymous):

it is not clear to me what this function is

OpenStudy (anonymous):

\[\ln(2)\cos(\frac{1}{4}-x)\]?

OpenStudy (anonymous):

\[\ln(2)\cos(\frac{1}{4-x})\]?

OpenStudy (anonymous):

no.... (ln(2))(cos( (1/4) -x ) // i dont understand y its not a negative number

OpenStudy (anonymous):

it is a function, not a number, so it could be either positive or negative

OpenStudy (anonymous):

the derivative of cosine is - sine

OpenStudy (anonymous):

but you are not taking the derivative of cosin(x) you are taking the derivative of cos (1/4 -x)

OpenStudy (anonymous):

so you have to use the chain rule and make sure to multply by the derivative of 1/4-x which is -1

OpenStudy (anonymous):

therefore you should get \[\ln(2)\sin(\frac{1}{4}-x)\] there is no minus sign becaues the derivative of \[\frac{1}{4}-x\] is -1

OpenStudy (anonymous):

ok..so heres how i see it. derivative of ln(2) cos( 1/4 -x ) = 0*cos( 1/4 -x) + ln(2) -sin ( 1/4 -x) .. right?

OpenStudy (anonymous):

ohhhh...... i forgot to multiply it by a (-1) for the derivative of the inside..right?

OpenStudy (anonymous):

yes you neede that one you do not need the product rule though

OpenStudy (anonymous):

\[\ln(2)\] is a constant

OpenStudy (anonymous):

ohhh.... i get it. thanks dude

OpenStudy (anonymous):

the derivative of cosine is - sine and the derivative of 4cosine is -4 sine don't use the product rule for constant multiples, it is a waste of time

OpenStudy (anonymous):

yw